OpenStudy (anonymous):

If sin (t) = 2/3, then what is csc (t)?

OpenStudy (luigi0210):

Welcome to openstudy, what you need to know is that: \[\sin t= \frac{opp}{hyp}\] and \[\csc t = \frac{hyp}{opp}\]